OpenStudy (anonymous):

next please: find the 1st, 4th and 10th terms of the arithmetic sequence described by the given rule: A(n)=12+(n-1)(3)

OpenStudy (anonymous):

I'm going to try and learn arithmetic sequences on my own, 7th grade doesn't have that.

OpenStudy (anonymous):

I thought you just substitute n with 1,4 and 10 to find your answer

OpenStudy (anonymous):

For 1 12+(1-1)(3)= 12 For 4 12+ (4-1)(3) = 21 For 10 12+ (10-1) (3) = 39

OpenStudy (anonymous):

how did you get the 1 4 and 10

OpenStudy (anonymous):

1st term 4th term and 10th term

OpenStudy (anonymous):

oh duh i am so dumb!so all i need to do is substitute it to the n's. OMG so embarrassing. thank you so much.

OpenStudy (anonymous):

No problem. =) I always forget stuff like that too haha

OpenStudy (anonymous):

thats how you do that? wow that is actually a simple thing, never done it before though.